# Regularization by regular noise: a numerical result

### v1

@@ 4. Strong convergence rate
$II_{1}$ $=\int_{s}^{t}(b-\mathcal{P}^{H}_{r-s_{1}}b)(B_{r}^{H}+\varphi_{r})\mathrm{d}r+\int_{s}^{t}\mathcal{P}^{H}_{r-s_{1}}b(B_{r}^{H}+\varphi_{r})-\mathcal{P}^{H}_{r-s_{1}}b({\mathbb{E}}^{s_{1}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r})\mathrm{d}r.$ (4.16)
@@
$\|II_{1}\|_{L^{p}_{\omega}}$ $\leqslant N\int_{s}^{t}\|\mathcal{P}^{H}_{r-s_{1}}b-b\|_{{\mathcal{C}}^{0}}+\|\mathcal{P}^{H}_{r-s_{1}}b\|_{{\mathcal{C}}^{\alpha}}(\|B_{r}^{H}-{\mathbb{E}}^{s_{1}}B^{H}_{r}\|^{\alpha}_{L^{\alpha p}_{\omega}}+\|\varphi_{r}-{\mathbb{E}}^{s_{1}}\varphi_{r}\|^{\alpha}_{L_{\omega}^{\alpha p}})\mathrm{d}r$
$\leqslant N\|b\|_{{\mathcal{C}}^{\alpha}}\int_{s}^{t}(r-s_{1})^{\alpha H}+(|r-s_{1}|^{\alpha H}+|r-s_{1}|^{\alpha(1+\alpha H)})\mathrm{d}r$
$\leqslant N\|b\|_{{\mathcal{C}}^{\alpha}}(t-s)^{(1+\alpha H)\wedge(1+\alpha(1+\alpha H))},$ (4.17)
@@
Finally, by Lemma˜3.1, the proof is complete. ∎
Let us move to estimate $\mathcal{E}^{b,n,2}$ term.
@@
Suppose Assumption˜2.1 holds. Then we have for any $p\geqslant 1$ and $(s,t)\in[0,1]_{0}^{2}$
$\left\|\mathcal{E}^{b,n,2}_{s,t}\right\|_{L^{p}_{\omega}}\leqslant\frac{N}{n}|t-s|^{\frac{1}{2}+\varepsilon}$ (4.18)
where $\epsilon>0$ is sufficiently small and $N=N(p,d,\alpha,H,\varepsilon,\|b\|_{{\mathcal{C}}^{\alpha}})$.
@@
$A_{s,t}:={\mathbb{E}}^{s-(t-s)}\int_{s}^{t}b(B_{r}^{H}+{\mathbb{E}}^{s-(t-s)}\varphi_{r}^{n})-b(B_{r}^{H}+{\mathbb{E}}^{s-(t-s)}\varphi_{k_{n}(r)}^{n})\mathrm{d}r.$ (4.19)
Similarly to (4) and (4.6), we have
$\|A_{s,t}\|_{L^{p}_{\omega}}$ $=\|\int_{s}^{t}\mathcal{P}^{H}_{r-[s-(t-s)]}b({\mathbb{E}}^{s-(t-s)}B^{H}_{r}+{\mathbb{E}}^{s-(t-s)}\varphi_{r}^{n})$
$\qquad-\mathcal{P}^{H}_{r-[s-(t-s)]}b({\mathbb{E}}^{s-(t-s)}B^{H}_{r}+{\mathbb{E}}^{s-(t-s)}\varphi^{n}_{k_{n}(r)})\mathrm{d}r\|_{L^{p}_{\omega}}$ (4.20)
$\leqslant N\|b\|_{{\mathcal{C}}^{\alpha}}\sup_{r\in[s,t]}\|\varphi_{r}^{n}-\varphi_{k_{n}(r)}^{n}\|_{L_{\omega}^{p}}|t-s|^{1-(1-\alpha)H}$
$\leqslant Nn^{-1}\|b\|_{{\mathcal{C}}^{\alpha}}|t-s|^{1-(1-\alpha)H},$ (4.21)
where in the second inequality we used $\|\varphi_{r}^{n}-\varphi_{k_{n}(r)}^{n}\|_{L_{\omega}^{p}}\leqslant\|b\|_{{\mathcal{C}}^{0}}n^{-1}$.
Then, with taking $C_{1}=N\|b\|_{{\mathcal{C}}^{\alpha}}n^{-1}$, (3.1) holds by the fact that $1-(1-\alpha)H>\frac{1}{2}.$
${\mathbb{E}}^{s_{1}}\delta A_{s,u,t}=$ ${\mathbb{E}}^{s_{1}}\int_{s_{4}}^{s_{5}}\mathcal{P}_{r-s_{3}}^{H}b({\mathbb{E}}^{s_{3}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r}^{n})-\mathcal{P}_{r-s_{3}}^{H}b({\mathbb{E}}^{s_{3}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n})$
$\qquad\qquad-\mathcal{P}^{H}_{r-s_{3}}b({\mathbb{E}}^{s_{3}}B_{r}^{H}+{\mathbb{E}}^{s_{3}}\varphi_{r}^{n})+\mathcal{P}_{r-s_{3}}^{H}b({\mathbb{E}}^{s_{3}}B_{r}^{H}+{\mathbb{E}}^{s_{3}}\varphi_{k_{n}(r)}^{n})\mathrm{d}r$
$+{\mathbb{E}}^{s_{1}}\int_{s_{5}}^{s_{6}}\mathcal{P}_{r-s_{2}}^{H}b({\mathbb{E}}^{s_{2}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r}^{n})-\mathcal{P}_{r-s_{2}}^{H}b({\mathbb{E}}^{s_{2}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n})$
$\qquad\qquad-\mathcal{P}_{r-s_{2}}^{H}b({\mathbb{E}}^{s_{2}}B_{r}^{H}+{\mathbb{E}}^{s_{2}}\varphi_{r}^{n})+\mathcal{P}_{r-s_{3}}^{H}b({\mathbb{E}}^{s_{2}}B_{r}^{H}+{\mathbb{E}}^{s_{2}}\varphi_{k_{n}(r)}^{n})\mathrm{d}r$
$=$ $:I_{1}+I_{2}.$
$\|I_{1}\|_{L_{\omega}^{p}}\leqslant$ $N\|b\|_{{\mathcal{C}}^{\alpha}}\int_{s_{4}}^{s_{5}}(r-s_{3})^{-(1-\alpha)H}\big\|{\mathbb{E}}^{s_{1}}|{\mathbb{E}}^{s_{3}}\varphi_{r}^{n}-{\mathbb{E}}^{s_{3}}\varphi_{k_{n}(r)}^{n}-{\mathbb{E}}^{s_{1}}\varphi_{r}^{n}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n}|\big\|_{L_{\omega}^{p}}$
$\qquad+(r-s_{3})^{-(2-\alpha)H}\big\||{\mathbb{E}}^{s_{1}}\varphi_{r}^{n}-{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n}|\cdot{\mathbb{E}}^{s_{1}}|{\mathbb{E}}^{s_{3}}\varphi_{k_{n}(r)}^{n}-{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n}|\big\|_{L_{\omega}^{p}}\mathrm{d}r.$ (4.22)
Similar to (4.10), we have
$\|{\mathbb{E}}^{s_{1}}|{\mathbb{E}}^{s_{3}}\varphi_{r}^{n}-{\mathbb{E}}^{s_{3}}\varphi_{k_{n}(r)}^{n}-{\mathbb{E}}^{s_{1}}\varphi_{r}^{n}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n}|\|_{L^{p}_{\omega}}\leqslant\|{\mathbb{E}}^{s_{1}}|(\varphi_{r}^{n}-\varphi_{k_{n}(r)}^{n})-{\mathbb{E}}^{s_{1}}(\varphi_{r}^{n}-\varphi_{k_{n}(r)}^{n})|\|_{L^{p}_{\omega}}.$ (4.23)
We note that
When $s_{1}\in[k_{n}(r),r],$ we have
[…]
$\leqslant$ $N(r-k_{n}(r))\|b\|_{{\mathcal{C}}^{\alpha}}{\mathbb{E}}^{s_{1}}(|B_{k_{n}(r)}^{H}-{\mathbb{E}}^{s_{1}}B_{k_{n}(r)}^{H}|^{\alpha}+|\varphi_{k_{n}(r)}^{n}-\varphi_{s_{1}}^{n}|^{\alpha}).$
Moreover, using (3.5) and (4.2), we have
$\leqslant$ $N(r-k_{n}(r))\|b\|_{{\mathcal{C}}^{\alpha}}(|k_{n}(r)-s_{1}|^{\alpha H}+|k_{n}(r)-s_{1}|^{\alpha})$
$\leqslant$ $\frac{N}{n}\|b\|_{{\mathcal{C}}^{\alpha}}|r-s_{1}|^{\alpha}$
where we used the fact $H>1$ in the second inequality. Plugging it into (4.23), we get
@@
$\|{\mathbb{E}}^{s_{1}}|{\mathbb{E}}^{s_{3}}\varphi_{r}^{n}-{\mathbb{E}}^{s_{3}}\varphi_{k_{n}(r)}^{n}-{\mathbb{E}}^{s_{1}}\varphi_{r}^{n}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n}|\|_{L^{p}_{\omega}}\leqslant\frac{N}{n}\|b\|_{{\mathcal{C}}^{\alpha}}|r-s_{1}|^{\alpha}.$ (4.24)
Meanwhile (4.9) and (4.2) yield
${\mathbb{E}}^{s_{1}}|{\mathbb{E}}^{s_{3}}\varphi_{k_{n}(r)}^{n}-{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n}|\leqslant N|k_{n}(r)-s_{1}|^{1+\alpha H}\leqslant N|r-s_{1}|^{1+\alpha H},$ (4.25)
$\|{\mathbb{E}}^{s_{1}}(\varphi_{r}^{n}-\varphi_{k_{n}(r)}^{n})\|_{L_{\omega}^{p}}\leqslant N\|\varphi_{\cdot}^{n}-\varphi_{k_{n}(\cdot)}^{n}\|_{C_{p}^{0}}\leqslant N\frac{\|b\|_{C^{0}}}{n}.$ (4.26)
Applying (4.24), (4.25) and (4.26) into (4.22) gives us
$\|I_{1}\|_{L^{p}_{\omega}}$ $\leqslant N\frac{\|b\|_{{\mathcal{C}}^{\alpha}}}{n}\int_{s_{4}}^{s_{5}}(r-s_{3})^{-(1-\alpha)H}(r-s_{1})^{\alpha}\mathrm{d}r+(r-s_{3})^{-(2-\alpha)H}(r-s_{1})^{1+\alpha H}\mathrm{d}r$
$\leqslant N\frac{\|b\|_{{\mathcal{C}}^{\alpha}}}{n}(t-s)^{(1+\alpha-(1-\alpha)H)\wedge(2+(2\alpha-2)H)}.$ (4.27)
@@
With the same bound on $I_{2}$ and Assumption˜2.1 implying $(1+\alpha-(1-\alpha)H)\wedge(2+(2\alpha-2)H)>1$, (3.2) holds with $C_{2}=\frac{N}{n}\|b\|_{{\mathcal{C}}^{\alpha}}$.
@@
Lastly we claim that the process $\mathcal{A}$ in (3.3) is indeed given by
$\mathcal{A}_{t}=\int_{0}^{t}b(B^{H}_{r}+\varphi_{r}^{n})-b(B^{H}_{r}+\varphi_{k_{n}(r)}^{n})\mathrm{d}r.$ (4.28)
To prove this claim, it suffices to show (3.4). Similarly to (4), we have
$\mathcal{A}t-\mathcal{A}_{s}-A_{s,t}$ $=\int_{s}^{t}b(B^{H}_{r}+\varphi_{r}^{n})-\mathcal{P}^{H}_{r-s_{1}}b({\mathbb{E}}^{s_{1}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi^{n}_{r})\mathrm{d}r$
$\quad-\int_{s}^{t}b(B_{r}^{H}+\varphi_{k_{n}(r)}^{n})-\mathcal{P}^{H}_{r-s_{1}}b({\mathbb{E}}^{s_{1}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n})\mathrm{d}r.$
Then the estimates can be done similarly to that in Lemma˜4.2.
Therefore, the claim holds by the uniqueness from Lemma˜3.1. Finally, by Lemma˜3.1, we can conclude (4.18). ∎
@@
Suppose Assumption˜2.1 holds. Then for any $p\geqslant 1$ and $(s,t)\in[0,1]_{0}^{2}$, we have
$\left\|\mathcal{E}^{b,n,3}_{s,t}\right\|_{L^{p}_{\omega}}\leqslant\frac{N}{n}|t-s|^{\frac{1}{2}+\varepsilon}$ (4.29)
with sufficiently small $\varepsilon>0$ and some constant $N=N(p,d,\alpha,H,\varepsilon,\|b\|_{{\mathcal{C}}^{\alpha}})$.
@@
$A_{s,t}={\mathbb{E}}^{s-(t-s)}\int_{s}^{t}b(B_{r}^{H}+{\mathbb{E}}^{s-(t-s)}\varphi_{k_{n}(r)}^{n})-b(B_{k_{n}(r)}^{H}+{\mathbb{E}}^{s-(t-s)}\varphi_{k_{n}(r)}^{n})\mathrm{d}r.$
@@
$\|A_{s,t}\|_{L^{p}_{\omega}}$ $\leqslant N\|b\|_{{\mathcal{C}}^{\alpha}}\int_{s}^{t}\big\||B^{H}_{r}-B^{H}_{k_{n}(r)}|^{\alpha}\big\|_{L^{p}_{\omega}}\mathrm{d}r\leqslant N\|b\|_{{\mathcal{C}}^{\alpha}}|t-s|\cdot\frac{1}{n^{\alpha}}$
$\leqslant N\|b\|_{{\mathcal{C}}^{\alpha}}|t-s|^{\frac{1}{2}+\varepsilon}\frac{1}{n^{\alpha+\frac{1}{2}-\varepsilon}}.$
For $H>1$ and $\alpha>1-\frac{1}{2H}$ we can take such sufficiently small $\varepsilon>0$ so that $\epsilon\in(0,\alpha+\frac{1}{2H}-1)$, which implies
$\|A_{s,t}\|_{L^{p}_{\omega}}\leqslant\frac{N}{n}\|b\|_{{\mathcal{C}}^{\alpha}}|t-s|^{\frac{1}{2}+\varepsilon}.$ (4.30)
@@
$A_{s,t}$ $=\int_{s_{4}}^{s_{6}}\mathcal{P}^{H}_{r-s_{1}}b({\mathbb{E}}^{s_{1}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n})-\mathcal{P}^{H}_{k_{n}(r)-s_{1}}b({\mathbb{E}}^{s_{1}}B_{k_{n}(r)}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n})\mathrm{d}r$
$=\int_{s_{4}}^{s_{6}}\mathcal{P}^{H}_{r-s_{1}}b({\mathbb{E}}^{s_{1}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n})-\mathcal{P}^{H}_{r-s_{1}}b({\mathbb{E}}^{s_{1}}B_{k_{n}(r)}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n})$
$\qquad+\mathcal{P}^{H}_{r-s_{1}}b({\mathbb{E}}^{s_{1}}B_{k_{n}(r)}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n})-\mathcal{P}^{H}_{k_{n}(r)-s_{1}}b({\mathbb{E}}^{s_{1}}B^{H}_{k_{n}(r)}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n})\mathrm{d}r$
@@
$\|IV_{1}\|_{L^{p}_{\omega}}$ $\leqslant\int_{s_{4}}^{s_{6}}\|\mathcal{P}^{H}_{r-s_{1}}b\|_{{\mathcal{C}}^{1}}\|{\mathbb{E}}^{s_{1}}(B_{r}^{H}-B_{k_{n}(r)}^{H})\|_{L^{p}_{\omega}}\mathrm{d}r$
$\leqslant\frac{N\|b\|_{{\mathcal{C}}^{\alpha}}}{n}\int_{s_{4}}^{s_{6}}(r-s_{1})^{-(1-\alpha)H}\mathrm{d}r$
$\leqslant\frac{N}{n}\|b\|_{{\mathcal{C}}^{\alpha}}|t-s|^{1-(1-\alpha)H}.$ (4.31)
@@
$\|\mathcal{P}^{H}_{t}b-\mathcal{P}_{s}^{H}b\|_{{\mathcal{C}}^{0}}\leqslant Ns^{-(2-\alpha)H}|t^{2H}-s^{2H}|\|b\|_{{\mathcal{C}}^{\alpha}},$
@@
$|IV_{2}|$ $\leqslant N\|b\|_{{\mathcal{C}}^{\alpha}}\int_{s_{4}}^{s_{6}}(k_{n}(r)-s_{1})^{-(2-\alpha)H}\left((r-s_{1})^{2H}-(k_{n}(r)-s_{1})^{2H}\right)\mathrm{d}r.$
@@
$\big||r-s_{1}|^{2H}-|k_{n}(r)-s_{1}|^{2H}\big|\leqslant N|r-k_{n}(r)||r-s_{1}|^{2H-1}\leqslant\frac{N}{n}|t-s|^{2H-1},$
@@
$|IV_{2}|\leqslant\frac{N}{n}\|b\|_{{\mathcal{C}}^{\alpha}}\int_{s_{4}}^{s_{6}}(t-s)^{2H-1-(2-\alpha)H}\mathrm{d}r=\frac{N}{n}\|b\|_{{\mathcal{C}}^{\alpha}}(t-s)^{\alpha H}.$ (4.32)
Then (4.30) togehter with (4.31) and (4.32) verifies (3.1) of Lemma˜3.1 with taking $C_{1}=\frac{N}{n}\|b\|_{{\mathcal{C}}^{\alpha}}$, since $((1-(1-\alpha)H)\wedge(\alpha H\big))>\frac{1}{2}$.
@@
$\qquad-\mathcal{P}_{r-s_{3}}^{H}b({\mathbb{E}}^{s_{3}}B^{H}_{r}+{\mathbb{E}}^{s_{3}}\varphi_{k_{n}(r)}^{n})+\mathcal{P}^{H}_{r-s_{3}}b({\mathbb{E}}^{s_{3}}B^{H}_{k_{n}(r)}+{\mathbb{E}}^{s_{3}}\varphi^{n}_{k_{n}(r)})\mathrm{d}r$
@@
$\qquad-\mathcal{P}^{H}_{r-s_{2}}b({\mathbb{E}}^{s_{2}}B^{H}_{r}+{\mathbb{E}}^{s_{2}}\varphi_{k_{n}(r)}^{n})+\mathcal{P}^{H}_{r-s_{2}}b({\mathbb{E}}^{s_{2}}B^{H}_{k_{n}(r)}+{\mathbb{E}}^{s_{2}}\varphi^{n}_{k_{n}(r)})\mathrm{d}r$
@@
$\leqslant N\int_{s_{4}}^{s_{5}}\|b\|_{{\mathcal{C}}^{\alpha}}(r-s_{3})^{-H(2-\alpha)}\|{\mathbb{E}}^{s_{1}}[|{\mathbb{E}}^{s_{3}}(B_{r}^{H}-B_{k_{n}(r)}^{H})|\cdot|{\mathbb{E}}^{s_{1}}\varphi^{n}_{k_{n}(r)}-{\mathbb{E}}^{s_{3}}\varphi^{n}_{k_{n}(r)}|]\|_{L^{p}_{\omega}}\mathrm{d}r.$ (4.33)
@@
$\leqslant N({\mathbb{E}}^{s_{1}}(B^{H}_{r}-B^{H}_{k_{n}(r)})^{2})^{\frac{1}{2}}\cdot(k_{n}(r)-s_{1})^{1+\alpha H}.$

### v2

@@ 4. Strong convergence rate
$II_{1}$ $=\int_{s}^{t}(f-\mathcal{P}^{H}_{r-s_{1}}f)(B_{r}^{H}+\varphi_{r})\mathrm{d}r+\int_{s}^{t}\mathcal{P}^{H}_{r-s_{1}}f(B_{r}^{H}+\varphi_{r})-\mathcal{P}^{H}_{r-s_{1}}f({\mathbb{E}}^{s_{1}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{r})\mathrm{d}r.$ (4.16)
@@
$\|II_{1}\|_{L^{p}_{\omega}}$ $\leqslant C\int_{s}^{t}\|\mathcal{P}^{H}_{r-s_{1}}f-f\|_{{\mathcal{C}}^{0}}+\|\mathcal{P}^{H}_{r-s_{1}}f\|_{{\mathcal{C}}^{\alpha}}(\|B_{r}^{H}-{\mathbb{E}}^{s_{1}}B^{H}_{r}\|^{\alpha}_{L^{\alpha p}_{\omega}}+\|\varphi_{r}-{\mathbb{E}}^{s_{1}}\varphi_{r}\|^{\alpha}_{L_{\omega}^{\alpha p}})\mathrm{d}r$
$\leqslant C\|f\|_{{\mathcal{C}}^{\alpha}}\int_{s}^{t}(r-s_{1})^{\alpha H}+(|r-s_{1}|^{\alpha H}+|r-s_{1}|^{\alpha(1+\alpha H)})\mathrm{d}r$
$\leqslant C\|f\|_{{\mathcal{C}}^{\alpha}}(t-s)^{(1+\alpha H)\wedge(1+\alpha(1+\alpha H))},$ (4.17)
@@
Finally, by Lemma˜3.1, the proof completes. ∎
Let us move to estimate $\mathcal{E}^{f,n,2}$ term.
@@
Suppose Assumption˜2.1 holds. Then we have for any $p\geqslant 1$, $\gamma\in(\frac{1}{2},1+(\alpha-1)H)$ and $(s,t)\in[0,1]_{0}^{2}$
$\left\|\mathcal{E}^{f,n,2}_{s,t}\right\|_{L^{p}_{\omega}}\leqslant\frac{C}{n}\|f\|_{{\mathcal{C}}^{\alpha}}|t-s|^{\gamma+\varepsilon},\quad(s,t)\in[0,1]_{0}^{2},f\in{\mathcal{C}}^{\alpha}$ (4.18)
where $\epsilon>0$ is sufficiently small and $C=C(p,d,\alpha,\gamma,H,\varepsilon,\|b\|_{{\mathcal{C}}^{\alpha}})$.
@@
$A_{s,t}:={\mathbb{E}}^{s-(t-s)}\int_{s}^{t}f(B_{r}^{H}+{\mathbb{E}}^{s-(t-s)}\varphi_{r}^{n})-f(B_{r}^{H}+{\mathbb{E}}^{s-(t-s)}\varphi_{k_{n}(r)}^{n})\mathrm{d}r.$ (4.19)
We may notice that the analysis of showing (4.6) and estimating (4.7) also work for (4.19). Therefore we omit the detailed proof here for showing the following estimates and present in Appendix˜A; in the end we get
$\|A_{s,t}\|_{L^{p}_{\omega}}$ $\leqslant C\|f\|_{{\mathcal{C}}^{\alpha}}\sup_{r\in[s,t]}\|\varphi_{r}^{n}-\varphi_{k_{n}(r)}^{n}\|_{L_{\omega}^{p}}|t-s|^{1-(1-\alpha)H}$
$\leqslant Cn^{-1}\|f\|_{{\mathcal{C}}^{\alpha}}|t-s|^{1-(1-\alpha)H},$ (4.20)
@@
$\|{\mathbb{E}}^{s_{1}}\delta A_{s,u,t}\|_{L^{p}_{\omega}}$ $\leqslant C\frac{\|f\|_{{\mathcal{C}}^{\alpha}}}{n}(t-s)^{(1+\alpha-(1-\alpha)H)\wedge(2+(2\alpha-2)H)},$ (4.21)
recall the definition of $s_{i},i=1,\dots,6$ in (4.3).
@@
Then, with taking $C_{1}=C\|f\|_{{\mathcal{C}}^{\alpha}}n^{-1}$, (3.1) holds by the fact that $1-(1-\alpha)H>\frac{1}{2}.$ Assumption˜2.1 implying $(1+\alpha-(1-\alpha)H)\wedge(2+(2\alpha-2)H)>1$, (3.2) holds with $C_{2}=\frac{C}{n}\|f\|_{{\mathcal{C}}^{\alpha}}$
@@
Similarly to Lemma˜4.2, we could verify that the process $\mathcal{A}$ in (3.3) is indeed given by
$\mathcal{A}_{t}=\int_{0}^{t}f(B^{H}_{r}+\varphi_{r}^{n})-f(B^{H}_{r}+\varphi_{k_{n}(r)}^{n})\mathrm{d}r.$ (4.22)
Therefore, the claim (4.18) holds by the uniqueness from Lemma˜3.1. ∎
@@
Suppose Assumption˜2.1 holds. Then for any $p\geqslant 1$, $\gamma\in(\frac{1}{2},1+(\alpha-1)H)$ and $(s,t)\in[0,1]_{0}^{2}$, we have
$\left\|\mathcal{E}^{f,n,3}_{s,t}\right\|_{L^{p}_{\omega}}\leqslant\frac{C}{n}\|f\|_{{\mathcal{C}}^{\alpha}}|t-s|^{\gamma+\varepsilon},\quad(s,t)\in[0,1]_{0}^{2},f\in{\mathcal{C}}^{\alpha}$ (4.23)
with sufficiently small $\varepsilon>0$ and some constant $C=C(p,d,\alpha,\gamma,H,\varepsilon,\|b\|_{{\mathcal{C}}^{\alpha}})$.
@@
$A_{s,t}={\mathbb{E}}^{s-(t-s)}\int_{s}^{t}f(B_{r}^{H}+{\mathbb{E}}^{s-(t-s)}\varphi_{k_{n}(r)}^{n})-f(B_{k_{n}(r)}^{H}+{\mathbb{E}}^{s-(t-s)}\varphi_{k_{n}(r)}^{n})\mathrm{d}r.$
@@
$\|A_{s,t}\|_{L^{p}_{\omega}}$ $\leqslant C\|f\|_{{\mathcal{C}}^{\alpha}}\int_{s}^{t}\big\||B^{H}_{r}-B^{H}_{k_{n}(r)}|^{\alpha}\big\|_{L^{p}_{\omega}}\mathrm{d}r\leqslant C\|f\|_{{\mathcal{C}}^{\alpha}}|t-s|\cdot\frac{1}{n^{\alpha}}$
$\leqslant C\|f\|_{{\mathcal{C}}^{\alpha}}|t-s|^{\gamma+\varepsilon}\frac{1}{n^{\alpha+1-\gamma-\varepsilon}}.$
Since $\alpha>\gamma$ we can take such sufficiently small $\varepsilon>0$ so that $\epsilon\in(0,\alpha-\gamma)$, which implies
$\|A_{s,t}\|_{L^{p}_{\omega}}\leqslant\frac{C}{n}\|f\|_{{\mathcal{C}}^{\alpha}}|t-s|^{\frac{1}{2}+\varepsilon}.$ (4.24)
@@
$A_{s,t}$ $=\int_{s_{4}}^{s_{6}}\mathcal{P}^{H}_{r-s_{1}}f({\mathbb{E}}^{s_{1}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n})-\mathcal{P}^{H}_{k_{n}(r)-s_{1}}f({\mathbb{E}}^{s_{1}}B_{k_{n}(r)}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n})\mathrm{d}r$
$=\int_{s_{4}}^{s_{6}}\mathcal{P}^{H}_{r-s_{1}}f({\mathbb{E}}^{s_{1}}B_{r}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n})-\mathcal{P}^{H}_{r-s_{1}}f({\mathbb{E}}^{s_{1}}B_{k_{n}(r)}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n})$
$\qquad+\mathcal{P}^{H}_{r-s_{1}}f({\mathbb{E}}^{s_{1}}B_{k_{n}(r)}^{H}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n})-\mathcal{P}^{H}_{k_{n}(r)-s_{1}}f({\mathbb{E}}^{s_{1}}B^{H}_{k_{n}(r)}+{\mathbb{E}}^{s_{1}}\varphi_{k_{n}(r)}^{n})\mathrm{d}r$
@@
$\|IV_{1}\|_{L^{p}_{\omega}}$ $\leqslant\int_{s_{4}}^{s_{6}}\|\mathcal{P}^{H}_{r-s_{1}}f\|_{{\mathcal{C}}^{1}}\|{\mathbb{E}}^{s_{1}}(B_{r}^{H}-B_{k_{n}(r)}^{H})\|_{L^{p}_{\omega}}\mathrm{d}r\leqslant\frac{C\|f\|_{{\mathcal{C}}^{\alpha}}}{n}\int_{s_{4}}^{s_{6}}(r-s_{1})^{-(1-\alpha)H}\mathrm{d}r$
$\leqslant\frac{C}{n}\|f\|_{{\mathcal{C}}^{\alpha}}|t-s|^{1-(1-\alpha)H}.$ (4.25)
@@
$\|\mathcal{P}^{H}_{t}f-\mathcal{P}_{s}^{H}f\|_{{\mathcal{C}}^{0}}\leqslant Cs^{-(2-\alpha)H}|t^{2H}-s^{2H}|\|f\|_{{\mathcal{C}}^{\alpha}},$
@@
$|IV_{2}|$ $\leqslant C\|f\|_{{\mathcal{C}}^{\alpha}}\int_{s_{4}}^{s_{6}}(k_{n}(r)-s_{1})^{-(2-\alpha)H}\left((r-s_{1})^{2H}-(k_{n}(r)-s_{1})^{2H}\right)\mathrm{d}r.$
@@
$\big||r-s_{1}|^{2H}-|k_{n}(r)-s_{1}|^{2H}\big|\leqslant C|r-k_{n}(r)||r-s_{1}|^{2H-1}\leqslant\frac{C}{n}|t-s|^{2H-1},$
@@
$|IV_{2}|\leqslant\frac{C}{n}\|f\|_{{\mathcal{C}}^{\alpha}}\int_{s_{4}}^{s_{6}}(t-s)^{2H-1-(2-\alpha)H}\mathrm{d}r=\frac{C}{n}\|f\|_{{\mathcal{C}}^{\alpha}}(t-s)^{\alpha H}.$ (4.26)
Then (4.24) togehter with (4.25) and (4.26) verifies (3.1) of Lemma˜3.1 with taking $C_{1}=\frac{C}{n}\|f\|_{{\mathcal{C}}^{\alpha}}$, since $((1-(1-\alpha)H)\wedge(\alpha H\big))>\gamma>\frac{1}{2}$.
@@
$\qquad-\mathcal{P}_{r-s_{3}}^{H}f({\mathbb{E}}^{s_{3}}B^{H}_{r}+{\mathbb{E}}^{s_{3}}\varphi_{k_{n}(r)}^{n})+\mathcal{P}^{H}_{r-s_{3}}f({\mathbb{E}}^{s_{3}}B^{H}_{k_{n}(r)}+{\mathbb{E}}^{s_{3}}\varphi^{n}_{k_{n}(r)})\mathrm{d}r$
@@
$\qquad-\mathcal{P}^{H}_{r-s_{2}}f({\mathbb{E}}^{s_{2}}B^{H}_{r}+{\mathbb{E}}^{s_{2}}\varphi_{k_{n}(r)}^{n})+\mathcal{P}^{H}_{r-s_{2}}f({\mathbb{E}}^{s_{2}}B^{H}_{k_{n}(r)}+{\mathbb{E}}^{s_{2}}\varphi^{n}_{k_{n}(r)})\mathrm{d}r$
@@
$\leqslant C\|f\|_{{\mathcal{C}}^{\alpha}}\int_{s_{4}}^{s_{5}}(r-s_{3})^{-H(2-\alpha)}\|{\mathbb{E}}^{s_{1}}[|{\mathbb{E}}^{s_{3}}(B_{r}^{H}-B_{k_{n}(r)}^{H})|\cdot|{\mathbb{E}}^{s_{1}}\varphi^{n}_{k_{n}(r)}-{\mathbb{E}}^{s_{3}}\varphi^{n}_{k_{n}(r)}|]\|_{L^{p}_{\omega}}\mathrm{d}r.$ (4.27)
@@
$\leqslant C({\mathbb{E}}^{s_{1}}(B^{H}_{r}-B^{H}_{k_{n}(r)})^{2})^{\frac{1}{2}}\cdot(k_{n}(r)-s_{1})^{1+\alpha H}.$